# Optimal low-rank approximations of posteriors for linear Gaussian inverse problems on Hilbert spaces

### v1

@@ B.2 Proofs of Section 4
To prove item (i), we first prove the forward correspondence and show that $\mathscr{C}_{r}^{-1}\subset\mathscr{P}_{r}$. Let $K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ such that $\mathscr{C}_{\textup{pr}}-KK^{*}\in\mathcal{C}_{r}$. This means that $\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$ and $\mathcal{C}_{\textup{pr}}-KK^{*}$ is injective and has nonnegative spectrum. The range condition $\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$ implies that $\mathcal{C}_{\textup{pr}}^{-1/2}K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ is a well-defined operator and $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{-1/2}K}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$. Because $\mathcal{C}_{\textup{pr}}^{-1/2}$ is injective, we have $\operatorname{rank}\left({\mathcal{C}_{\textup{pr}}^{-1/2}K}\right)=\operatorname{rank}\left({K}\right)$. Since $K^{*},\mathcal{C}_{\textup{pr}}^{-1/2}$ and $K^{*}\mathcal{C}_{\textup{pr}}^{-1/2}$ are all densely defined, and since $\mathcal{C}_{\textup{pr}}^{-1/2}$ is self adjoint, by Lemma A.13(i), with $A\leftarrow\mathcal{C}_{\textup{pr}}^{-1/2}$, and $B\leftarrow K$ that $\mathcal{C}_{\textup{pr}}^{-1/2}KK^{*}\mathcal{C}_{\textup{pr}}^{-1/2}\subset\mathcal{C}_{\textup{pr}}^{-1/2}K(\mathcal{C}_{\textup{pr}}^{-1/2}K)^{*}$. Since $\mathcal{C}_{\textup{pr}}^{-1/2}K(\mathcal{C}_{\textup{pr}}^{-1/2}K)^{*}$ has the same rank as $K$ by Lemma A.8, and is self-adjoint and nonnegative, it follows that $\mathcal{C}_{\textup{pr}}^{-1/2}KK^{*}\mathcal{C}_{\textup{pr}}^{-1/2}$ can be extended by a diagonalisable operator, say, by $\sum_{i=1}^{\operatorname{rank}\left({K}\right)}d_{i}^{2}e_{i}\otimes e_{i}$, for some orthonormal sequence $(e_{i})_{i}\subset\mathcal{H}$ and sequence $(d_{i}^{2})_{i}$ with $d_{i}^{2}\geq d_{i+1}^{2}>0$. Now, $\mathcal{C}_{\textup{pr}}^{-1/2}KK^{*}\mathcal{C}_{\textup{pr}}^{-1/2}(\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}})\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$, hence $\mathcal{C}_{\textup{pr}}^{-1/2}K(\mathcal{C}_{\textup{pr}}^{-1/2}K)^{*}$ maps $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ into itself. Furthermore, $\mathcal{C}_{\textup{pr}}^{-1/2}K(\mathcal{C}_{\textup{pr}}^{-1/2}K)^{*}$ has closed range because it has finite rank. By density of $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}\subset\mathcal{H}$ and Lemma A.1, it follows that $\operatorname{ran}{\sum_{i=1}^{\operatorname{rank}\left({K}\right)}d_{i}^{2}e_{i}\otimes e_{i}}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}.$ Thus, $(e_{i})_{i}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ for $i\leq\operatorname{rank}\left({K}\right)$. Since $\mathcal{C}\in\mathscr{C}_{r}$, we then see that
$\mathcal{C}_{\textup{pr}}-KK^{*}=\mathcal{C}_{\textup{pr}}^{1/2}(I-\mathcal{C}_{\textup{pr}}^{-1/2}KK^{*}\mathcal{C}_{\textup{pr}}^{-1/2})\mathcal{C}_{\textup{pr}}^{1/2}=\mathcal{C}_{\textup{pr}}^{1/2}(I-\sum_{i=1}^{\operatorname{rank}\left({K}\right)}d_{i}^{2}e_{i}\otimes e_{i})\mathcal{C}_{\textup{pr}}^{1/2},$ (36)
is injective and positive. If $d_{i}^{2}=1$ for some $i\leq\operatorname{rank}\left({K}\right)$, then $\mathcal{C}_{\textup{pr}}^{-1/2}e_{i}\in\ker{(\mathcal{C}_{\textup{pr}}-KK^{*})}$, contradicting injectivity. If $d_{i}^{2}>1$ for some $i\leq\operatorname{rank}\left({K}\right)$, then $\langle(\mathcal{C}_{\textup{pr}}-KK^{*})\mathcal{C}_{\textup{pr}}^{-1/2}e_{i},\mathcal{C}_{\textup{pr}}^{-1/2}e_{i}\rangle<0$, contradicting positivity. Thus, $d_{i}^{2}\in(0,1)$ for all $i\leq\operatorname{rank}\left({K}\right)$. Hence, $I-\sum_{i=1}^{\operatorname{rank}\left({K}\right)}d_{i}^{2}e_{i}\otimes e_{i}$ maps $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ onto itself. It follows from (36) that
$\operatorname{ran}{\mathcal{C}_{\textup{pr}}-KK^{*}}=\mathcal{C}_{\textup{pr}}^{1/2}\left(I-\sum_{i=1}^{\operatorname{rank}\left({K}\right)}d_{i}^{2}e_{i}\otimes e_{i}\right)(\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}})=\mathcal{C}_{\textup{pr}}^{1/2}(\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}})=\operatorname{ran}{\mathcal{C}_{\textup{pr}}}.$
Because $(-d_{i}^{2})_{i}\in\ell^{2}((-1,\infty))$, we can then write the inverse that $\mathcal{C}_{\textup{pr}}-KK^{*}$ has on $\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$ as, using (36) and Lemma A.23 with $\delta_{i}\leftarrow(-d_{i}^{2})$
$(\mathcal{C}_{\textup{pr}}-KK^{*})^{-1}$ $=\mathcal{C}_{\textup{pr}}^{-1/2}(I-\sum_{i=1}^{\operatorname{rank}\left({K}\right)}d_{i}^{2}e_{i}\otimes e_{i})^{-1}\mathcal{C}_{\textup{pr}}^{-1/2}$
$=\mathcal{C}_{\textup{pr}}^{-1/2}(I+\sum_{i=1}^{\operatorname{rank}\left({K}\right)}\frac{d_{i}^{2}}{1-d_{i}^{2}}e_{i}\otimes e_{i})\mathcal{C}_{\textup{pr}}^{-1/2}$
$=\mathcal{C}_{\textup{pr}}^{-1}+UU^{*},$
where
$U:=\mathcal{C}_{\textup{pr}}^{-1/2}\sum_{i=1}^{\operatorname{rank}\left({K}\right)}\frac{d_{i}}{\sqrt{1-d_{i}^{2}}}e_{i}\otimes\phi_{i}\in\mathcal{B}(\mathbb{R}^{r},H),$
for any choice of ONB $(\phi_{i})_{i=1}^{r}$ of $\mathbb{R}^{r}$. This proves that $\mathscr{C}_{r}^{-1}\subset\mathscr{P}_{r}$. Since $\mathcal{C}_{\textup{pr}}^{-1/2}$ is injective, $\operatorname{rank}\left({U}\right)=\operatorname{rank}\left({K}\right)$ and this proves the forward correspondence states in item (i).
Next, we prove $\mathscr{C}_{r}^{-1}\supset\mathscr{P}_{r}$. Let $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. Then $\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}\in\mathscr{P}_{r}$. The nonnegative, self-ajoint and finite rank operator $\mathcal{C}_{\textup{pr}}^{1/2}UU^{*}\mathcal{C}_{\textup{pr}}^{1/2}$ has the same rank as $\mathcal{C}_{\textup{pr}}^{1/2}UU^{*}$ by Lemma A.1 and injectivity of $\mathcal{C}_{\textup{pr}}^{1/2}$. In turn, $\mathcal{C}_{\textup{pr}}^{1/2}UU^{*}$ has the same rank as $U$. We can therefore diagonalise $\mathcal{C}_{\textup{pr}}^{1/2}UU^{*}\mathcal{C}_{\textup{pr}}^{1/2}$ as, say, $\sum_{i=1}^{\operatorname{rank}\left({U}\right)}c_{i}^{2}f_{i}\otimes f_{i}$ with the $(f_{i})_{i}$ orthonormal, $(f_{i})_{i}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ for $i\leq\operatorname{rank}\left({U}\right)$, and $c_{i}^{2}\geq c_{i+i}^{2}>0$. Thus, $I+\sum_{i=1}^{\operatorname{rank}\left({U}\right)}c_{i}^{2}f_{i}\otimes f_{i}\geq I$. In particular, $I+\sum_{i=1}^{\operatorname{rank}\left({U}\right)}c_{i}^{2}f_{i}\otimes f_{i}$ is a bijection, and maps $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ onto itself. Hence, as a map $\operatorname{ran}{\mathcal{C}_{\textup{pr}}}\rightarrow\mathcal{H}$, it holds that
$\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}=\mathcal{C}_{\textup{pr}}^{-1/2}(I+\mathcal{C}_{\textup{pr}}^{1/2}UU^{*}\mathcal{C}_{\textup{pr}}^{1/2})\mathcal{C}_{\textup{pr}}^{-1/2}=\mathcal{C}_{\textup{pr}}^{-1/2}\left(I+\sum_{i=1}^{\operatorname{rank}\left({U}\right)}c_{i}^{2}f_{i}\otimes f_{i}\right)\mathcal{C}_{\textup{pr}}^{-1/2},$ (37)
is a bijection, since it is a composition of bijections. Since $UU^{*}\in\mathcal{B}(\mathcal{H})$, $\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}$ is closed by Lemma A.10. As a closed linear bijection, $\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}$ has a closed and everywhere defined inverse by Lemma A.10, which is then bounded by Lemma A.11. By (37), by the fact that $(c_{i})_{i}\in\ell^{2}((-1,\infty)$) and by Lemma A.23 with $\delta_{i}\leftarrow c_{i}^{2}$, this inverse equals,
$\left(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}\right)^{-1}$ $=\mathcal{C}_{\textup{pr}}^{1/2}(I+\sum_{i=1}^{\operatorname{rank}\left({U}\right)}c_{i}^{2}f_{i}\otimes f_{i})^{-1}\mathcal{C}_{\textup{pr}}^{1/2}$
$=\mathcal{C}_{\textup{pr}}^{1/2}(I-\sum_{i=1}^{\operatorname{rank}\left({U}\right)}\frac{c_{i}^{2}}{1+c_{i}^{2}}f_{i}\otimes f_{i})^{-1}\mathcal{C}_{\textup{pr}}^{1/2}$
$=\mathcal{C}_{\textup{pr}}-KK^{*},$
where
$K\coloneqq\mathcal{C}_{\textup{pr}}^{1/2}\sum_{i=1}^{\operatorname{rank}\left({U}\right)}\frac{c_{i}}{\sqrt{1+c_{i}^{2}}}f_{i}\otimes\phi_{i}\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H}),$
for any choice of ONB $(\phi_{i}^{\prime})_{i=1}^{r}$ of $\mathbb{R}^{r}$. Since $(f_{i})_{i=1}^{\operatorname{rank}\left({U}\right)}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$, we have $\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$. We now show that $\mathcal{C}_{\textup{pr}}-KK^{*}>0$. Denote by $\sigma_{p}(A)$ the point spectrum of an operator $A$, that is, the set of eigenvalues. Then by Lemma A.18, $\sigma_{p}(\mathcal{C}_{\textup{pr}}^{-1})=\sigma_{p}(\mathcal{C}_{\textup{pr}})^{-1}$ and $\sigma_{p}(\mathcal{C}_{\textup{pr}}-KK^{*})=\sigma_{p}(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1}$. Let $\lambda$ be an eigenvalue of $\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}$ with eigenvector $v$. If $v\in\ker{U^{*}}$, then $\lambda\in\sigma_{p}(\mathcal{C}_{\textup{pr}}^{-1})=\sigma_{p}(\mathcal{C}_{\textup{pr}})^{-1}$ and thus $\lambda>0$ because $\mathcal{C}_{\textup{pr}}>0$. If $v\not\in\ker{U^{*}}$, then $\lambda\lVert v\rVert^{2}=\langle(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})v,v\rangle\geq\lVert U^{*}v\rVert^{2}>0$. Thus $\sigma_{p}(\mathcal{C}_{\textup{pr}}-KK^{*})=\sigma_{p}(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1}\subset(0,\infty)$. It follows that $\sigma(\mathcal{C}_{\textup{pr}}-KK^{*})=\left\{0\right\}\cup\sigma_{p}(\mathcal{C}_{\textup{pr}}-KK^{*})\subset[0,\infty)$. We thus conclude that $(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1}\in\mathscr{C}_{r}$. This implies that $\mathscr{P}_{r}^{-1}\subset\mathscr{C}_{r}$ and hence $\mathscr{P}_{r}\supset\mathscr{C}_{r}$. Finally, $\operatorname{rank}\left({K}\right)=\operatorname{rank}\left({U}\right)$ by injectivity of $\mathcal{C}_{\textup{pr}}^{1/2}$, thus proving the reverse corresponcence stated in the lemma. ∎
See 4.6
@@
The second statement follows from the first statement, since $\mathcal{L}\in\mathscr{L}$ is defined on $\mathcal{E}^{2}$ by definition (15b), and since $\mathscr{P}_{r}^{-1}=\mathscr{C}_{r}$ by Lemma 4.5(i). Let $\mathcal{C}\in\mathscr{C}_{r}$ arbitrary. Thus $\mathcal{C}=\mathcal{C}_{\textup{pr}}-KK^{*}$ for $K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ and $\mathcal{C}>0$ Since both $\mathcal{C}_{\textup{pr}}$ and $KK^{*}$ belong to $L_{1}(\mathcal{H})_{\mathbb{R}}$, it follows that $\mathcal{C}\in L_{1}(\mathcal{H})_{\mathbb{R}}$. Thus, by (8), it remains to show that $\mathcal{C}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}$ satisfies property E, as defined in Definition 3.2. As the set of operators which satisfy Property E is closed under taking adjoints, it suffices to show that $(\mathcal{C}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2})^{*}$ satisfies property E. By Lemma A.20 and Definition 3.2, we must then show $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}=\operatorname{ran}{\mathcal{C}^{1/2}}$, and $(\mathcal{C}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2})(\mathcal{C}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2})^{*}-I\in L_{2}(\mathcal{H})$.
By Lemma 4.5(i), $\mathcal{C}^{-1}=\mathcal{C}_{0}^{-1}+UU^{*}$ for some $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ with $\operatorname{rank}\left({U}\right)=\operatorname{rank}\left({K}\right)$. Since $\operatorname{dom}{UU^{*}}=\mathcal{H}$,
$\operatorname{ran}{\mathcal{C}}=\operatorname{dom}{\mathcal{C}^{-1}}=\operatorname{dom}{\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}}=\operatorname{dom}{\mathcal{C}_{\textup{pr}}^{-1}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}}.$
Let $\sum_{i=1}^{\operatorname{rank}\left({K}\right)}c_{i}f_{i}\otimes f_{i}$ be a diagonalisation of the self-adjoint, $\operatorname{rank}\left({K}\right)$-operator $\mathcal{C}_{\textup{pr}}^{1/2}UU^{*}\mathcal{C}_{\textup{pr}}^{1/2}$, with $(f_{i})_{i}$ an ONB of $\mathcal{H}$, $f_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ for all $i\leq\operatorname{rank}\left({K}\right)$ and $c_{i}^{2}\geq c_{i+1}^{2}\geq 0$. On $\operatorname{ran}{\mathcal{C}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$ it thus holds that
$\mathcal{C}^{-1}=\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}=\mathcal{C}_{\textup{pr}}^{-1/2}(I+\mathcal{C}_{\textup{pr}}^{1/2}UU^{*}\mathcal{C}_{\textup{pr}}^{1/2})\mathcal{C}_{\textup{pr}}^{-1/2}=\mathcal{C}_{\textup{pr}}^{-1/2}(I+\sum_{i=1}^{\operatorname{rank}\left({K}\right)}c_{i}^{2}f_{i}\otimes f_{i})\mathcal{C}_{\textup{pr}}^{-1/2}.$ (38)
By Lemma A.13(i), we have $(\mathcal{C}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2})^{*}=\mathcal{C}_{\textup{pr}}^{1/2}\mathcal{C}^{-1/2}$ on $\operatorname{dom}{\mathcal{C}^{-1/2}}=\operatorname{ran}{\mathcal{C}^{1/2}}\subset\operatorname{ran}{\mathcal{C}}$. Therefore, on $\operatorname{ran}{\mathcal{C}}=\operatorname{ran}{\mathcal{C}}_{\textup{pr}}$
$(\mathcal{C}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2})^{*}(\mathcal{C}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2})-I=\mathcal{C}_{\textup{pr}}^{1/2}\mathcal{C}^{-1}\mathcal{C}_{\textup{pr}}^{1/2}-I=\sum_{i=1}^{\operatorname{rank}\left({K}\right)}c_{i}^{2}f_{i}\otimes f_{i}.$
Since $\operatorname{ran}{\mathcal{C}_{\textup{pr}}}\subset\mathcal{H}$ is dense, this equality holds on all of $\mathcal{H}$, whence $(\mathcal{C}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2})^{*}(\mathcal{C}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2})-I\in L_{2}(\mathcal{H})$. We now show that $\operatorname{ran}{\mathcal{C}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$. Recall the definion $\lVert\cdot\rVert_{\mathcal{C}_{\textup{pr}}^{-1}}$ of the Cameron–Martin norm on $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ as $\lVert h\rVert_{\mathcal{C}_{\textup{pr}}^{-1}}=\lVert\mathcal{C}_{\textup{pr}}^{-1/2}\rVert$ from Section 2. Similarly, we define the norm $\lVert\cdot\rVert_{\mathcal{C}^{-1}}$ on $\operatorname{ran}{\mathcal{C}^{1/2}}$. The Cameron–Martin space $(\operatorname{ran}{C}_{\textup{pr}}^{1/2},\lVert\cdot\rVert_{\mathcal{C}_{\textup{pr}}^{-1}})$ can equivalently be defined as a completion of $\operatorname{ran}{\mathcal{C}}_{\textup{pr}}$ with respect to $\lVert\cdot\rVert_{\mathcal{C}_{\textup{pr}}^{-1}}$, see [16, Definition 3.49]. The analogous statement holds for $(\operatorname{ran}{C}^{1/2},\lVert\cdot\rVert_{\mathcal{C}^{-1}})$. We can thus show $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}$ holds by showing that the norms $\lVert\cdot\rVert_{\mathcal{C}^{-1}}$ and $\lVert\cdot\rVert_{\mathcal{C}_{\textup{pr}}^{-1}}$ are equivalent on $\operatorname{ran}{\mathcal{C}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$, as it then follows that also the completions of $\operatorname{ran}{\mathcal{C}}$ and $\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$ in either of the equivalent Cameron–Martin norms agree.
Now, on the one hand, $\langle\mathcal{C}^{-1}h,h\rangle=\langle(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})h,h\rangle\geq\langle\mathcal{C}_{\textup{pr}}^{-1}h,h\rangle$ for $h\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$. On the other hand, for any $h\in\operatorname{ran}{C}_{\textup{pr}}$ we have by (38) and the fact that $\mathcal{C}_{\textup{pr}}^{-1/2}$ is self-adjoint
$\langle\mathcal{C}^{-1}h,h\rangle$ $=\langle\mathcal{C}_{\textup{pr}}^{-1/2}(I+\sum_{i=1}^{\operatorname{rank}\left({K}\right)}c_{i}^{2}f_{i}\otimes f_{i})\mathcal{C}_{\textup{pr}}^{-1/2}h,h\rangle$
$=\langle(I+\sum_{i=1}^{\operatorname{rank}\left({K}\right)}c_{i}^{2}f_{i}\otimes f_{i})\mathcal{C}_{\textup{pr}}^{-1/2}h,\mathcal{C}_{\textup{pr}}^{-1/2}h\rangle$
$=\lVert\mathcal{C}_{\textup{pr}}^{-1/2}h\rVert^{2}+\sum_{i=1}^{\operatorname{rank}\left({K}\right)}c_{i}^{2}\langle f_{i},\mathcal{C}_{\textup{pr}}^{-1/2}h\rangle^{2}.$
Because and $(c_{i}^{2})_{i}$ is nonincreasing, this implies
$\langle\mathcal{C}^{-1}h,h\rangle$ $\leq\lVert\mathcal{C}_{\textup{pr}}^{-1/2}h\rVert^{2}+c_{1}^{2}\sum_{i=1}^{\operatorname{rank}\left({K}\right)}\langle f_{i},\mathcal{C}_{\textup{pr}}^{-1/2}h\rangle^{2}$
$\leq\lVert\mathcal{C}_{\textup{pr}}^{-1/2}h\rVert^{2}+c_{1}^{2}\sum_{i=1}^{\infty}\langle f_{i},\mathcal{C}_{\textup{pr}}^{-1/2}h\rangle^{2}$
$\leq(1+c_{1}^{2})\langle\mathcal{C}_{\textup{pr}}^{-1}h,h\rangle.$
Thus $\operatorname{ran}{\mathcal{C}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$, which concludes the proof.
See 4.7
Proof of Lemma 4.7.
To see that the self-adjoint operator $g(U)$ is a Hilbert–Schmidt operator, notice that by Lemma 4.6, $(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1}\in\mathcal{E}$ for $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. That is, by Lemma 3.3(ii) applied with $\mathcal{C}_{1}\leftarrow(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1}$ and $\mathcal{C}_{2}\leftarrow\mathcal{C}_{\textup{pos}}$, $g(U)$ is Hilbert–Schmidt and
$\mathcal{C}_{\textup{pos}}^{1/2}(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})\mathcal{C}_{\textup{pos}}^{1/2}-I\subset g(U)=\sum_{i}\gamma_{i}e_{i}\otimes e_{i},$
@@
for some ONB $(e_{i})_{i}$ of $\mathcal{H}$ and eigenvalue sequence $(\gamma_{i})_{i}\in\ell^{2}((-1,\infty))$. By Lemma 3.3(i)-(ii) with $\mathcal{C}_{1}\leftarrow\mathcal{C}_{\textup{pr}}$ and $\mathcal{C}_{2}\leftarrow\mathcal{C}_{\textup{pos}}$, and by Lemma 3.6, we have that $\operatorname{rank}\left({(\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2})^{*}\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2}-I}\right)=\operatorname{rank}\left({H}\right)$ and
@@
$\subset(\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2})^{*}\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2}-I+\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}.$
As $\mathcal{C}_{\textup{pos}}^{1/2}(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})\mathcal{C}_{\textup{pos}}^{1/2}-I$ is continuous and densely defined, the above extension is unique, which shows that (18) holds. Now, by injectivity of $\mathcal{C}_{\textup{pos}}^{1/2}$, we conclude from (18) that
$\operatorname{rank}\left({g(U)}\right)\leq\operatorname{rank}\left({(\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2})^{*}\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2}-I)}\right)+\operatorname{rank}\left({\mathcal{C}_{\textup{pos}}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}}\right)=\operatorname{rank}\left({H}\right)+r.$
and that $\operatorname{ran}{g(U)}\subset\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$. For $i\leq\operatorname{rank}\left({H}\right)+r$ we have $\gamma_{i}\not=0$ and it follows that $e_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ for $i\leq\operatorname{rank}\left({H}\right)+r$. By Lemma A.2, we can extend $(e_{i})_{i=1}^{\operatorname{rank}\left({H}\right)+r}$ to a sequence in $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ which is an ONB of $\mathcal{H}$. By replacing $(e_{i})_{i}$ with this sequence, we still have $g(U)=\sum_{i}\gamma_{i}e_{i}\otimes e_{i}$ and now $e_{i}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ for all $i$.
In (17), the order of $(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{1/2}\mathcal{C}_{\textup{pos}}^{1/2}$ and $\bigr{(}(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{1/2}\mathcal{C}_{\textup{pos}}^{1/2}\bigr{)}^{*}$ is reversed. Therefore, by comparing Lemma 3.3 (i) and (ii) with the same choices stated above, it follows that the eigenvalues of $R(\mathcal{C}_{\textup{pos}}\|(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1})$ and $g(U)$ agree, counting multiplicities. ∎
See 4.9

### v3

@@ B.2 Proofs of Section 4
Item (i) follows directly from Proposition 4.6(iii) applied with $(m_{1},\mathcal{C}_{1})\leftarrow(0,\mathcal{C}_{\textup{pr}})$ and the definitions (4) and (5).
See 4.9
Proof of Corollary 4.9.
The second statement follows from the first statement, since $\mathscr{P}_{r}^{-1}=\mathscr{C}_{r}$ by Corollary 4.8(i), and since $\mathcal{L}\in\mathscr{L}$ is finite on $\mathcal{E}^{2}$ by definition (16b) and by Lemma 4.1(i). The first statement follows from Proposition 4.6(ii)-(iii) applied with $(m_{1},\mathcal{C}_{1})\leftarrow(0,\mathcal{C}_{\textup{pr}})$ and the definition (4) of $\mathscr{C}_{r}$.
See 4.10
@@
By Corollary 4.8(i) and Corollary 4.9, $(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1}\in\mathcal{E}$ for $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. By Lemma 3.4(ii) applied with $\mathcal{C}_{1}\leftarrow(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1}$ and $\mathcal{C}_{2}\leftarrow\mathcal{C}_{\textup{pos}}$, $\mathcal{C}_{\textup{pos}}^{1/2}(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})\mathcal{C}_{\textup{pos}}^{1/2}-I$ is densely defined and there exists an ONB $(e_{i})_{i}$ of $\mathcal{H}$ and eigenvalue sequence $(\gamma_{i})_{i}\in\ell^{2}((-1,\infty))$ such that
$\mathcal{C}_{\textup{pos}}^{1/2}(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})\mathcal{C}_{\textup{pos}}^{1/2}-I\subset((\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{1/2}\mathcal{C}_{\textup{pos}}^{1/2})^{*}((\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{1/2}\mathcal{C}_{\textup{pos}}^{1/2})-I=\sum_{i}\gamma_{i}e_{i}\otimes e_{i},$
and by comparing with the expansion in Lemma 3.4(i), $\sum_{i}\gamma_{i}e_{i}\otimes e_{i}$ has the same eigenvalues as $R(\mathcal{C}_{\textup{pos}}\|(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1})$, counting multiplicities. Using (10b), the leftmost operator can be written as
@@
$\subset(\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2})^{*}\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2}-I+\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}$
$=\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}-\mathcal{C}_{\textup{pos}}^{1/2}H\mathcal{C}_{\textup{pos}}^{1/2}=g(U).$
Since $\mathcal{C}_{\textup{pos}}^{1/2}(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})\mathcal{C}_{\textup{pos}}^{1/2}-I$ is densely defined, the above continuous extension is unique, which shows that $g(U)=\sum_{i}\gamma_{i}e_{i}\otimes e_{i}$. By Proposition 3.7, $\operatorname{rank}\left({\mathcal{C}_{\textup{pos}}^{1/2}H\mathcal{C}_{\textup{pos}}^{1/2}}\right)=\operatorname{rank}\left({H}\right)$. Thus,
$\operatorname{rank}\left({g(U)}\right)\leq\operatorname{rank}\left({\mathcal{C}_{\textup{pos}}^{1/2}H\mathcal{C}_{\textup{pos}}^{1/2}}\right)+\operatorname{rank}\left({\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}}\right)\leq\operatorname{rank}\left({H}\right)+r.$
Furthermore, $\operatorname{ran}{g(U)}\subset\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$. For $i\in\mathbb{N}$ such that $\gamma_{i}\not=0$, this implies $e_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$. By Lemma A.1, we can extend $(e_{i})_{i:\gamma_{i}\not=0}$ to a sequence in $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ which is an ONB of $\mathcal{H}$. Replacing $(e_{i})_{i}$ with this sequence, we still have $g(U)=\sum_{i}\gamma_{i}e_{i}\otimes e_{i}$ and now $e_{i}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ for all $i$. ∎
See 4.11